# ITP-CAS-25-215

### v1

@@
Hai-Jun Lia,b, Wei Chaoc,d, Huai-Ke Guob,e, and Yu-Feng Zhoua,b,e,f
aInstitute of Theoretical Physics, Chinese Academy of Sciences, Beijing 100190, China
bInternational Centre for Theoretical Physics Asia-Pacific, Beijing 100190, China
cCenter for Advanced Quantum Studies, School of Physics and Astronomy, Beijing Normal University, Beijing 100875, China
dKey Laboratory of Multi-scale Spin Physics, Ministry of Education, Beijing Normal University, Beijing 100875, China
eSchool of Physical Sciences, University of Chinese Academy of Sciences, Beijing 100049, China
fSchool of Fundamental Physics and Mathematical Sciences, Hangzhou Institute for Advanced Study, UCAS, Hangzhou 310024, China
@@
In this work, we present the refinement of axion double level crossings within the context of multi-axion mass mixing, specifically focusing on cases where the number of axions exceeds two. Our investigation reveals that double level crossings are a common phenomenon in the mass mixing of the $Z_{\mathcal{N}}$ axion and axion-like particles. We introduce the general model for double level crossings, along with several toy examples, and redefine the light and heavy axion scenarios. In the light axion scenario, double level crossings can occur multiple times in the large $\mathcal{N}$} limit. However, excessively large values of $\mathcal{N}$} may also prevent the occurrence of double level crossings. Conversely, in the heavy axion scenario, excessively small $\mathcal{N}$} may similarly prevent their occurrence. Our findings also have some intriguing implications for axion cosmology.
August 18, 2025
@@ 1 Introduction
Within such a theoretical framework of the axiverse, it is natural to consider the cosmological evolution of multiple axions, a subject that has garnered considerable attention in recent times Hill:1988bu ; Daido:2015cba ; Cyncynates:2023esj ; Li:2023uvt ; Li:2025cep . When considering non-zero mass mixing between axions, a significant evolution known as level crossing can occur during the QCD phase transition. Typically, level crossing precedes the critical temperature of the QCD phase transition, denoted as $T_{\rm QCD}$. However, as emphasized in ref. Li:2023uvt , this rule is not absolute; in scenarios featuring a $Z_{\mathcal{N}}$ axion, [^1] In the $Z_{\mathcal{N}}$ axion scenario Hook:2018jle ; DiLuzio:2021pxd , $\mathcal{N}$} mirror and degenerate worlds, nonlinearly realized through the axion field under a $Z_{\mathcal{N}}$ symmetry, can potentially coexist in Nature. Notably, one of these worlds corresponds to the SM world. To address the strong CP problem in this context, it is imperative that $\mathcal{N}$} be an odd integer satisfying ${\mathcal{N}}\geqslant 3$. The $Z_{\mathcal{N}}$ axion can also serve as a candidate for DM DiLuzio:2021gos . a second level crossing might occur precisely at $T_{\rm QCD}$, a phenomenon termed double level crossings (see also ref. Li:2024kdy for another illustration). Axion level crossing can induce the adiabatic or non-adiabatic transitions in axion energy density, similar to the MSW effect Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj in neutrino oscillations, and has some profound cosmological implications Kitajima:2014xla ; Daido:2015bva ; Ho:2018qur ; Cyncynates:2021xzw ; Cyncynates:2022wlq ; Li:2023xkn ; Li:2024psa ; Murai:2024nsp .
@@
The rest of this paper is structured as follows. In section 2, we present the general model for axion double level crossings. In section 3, we present several toy examples of double level crossings in the light axion scenario. Then, we redefine the light and heavy axion scenarios in section 4. Finally, the conclusion is given in section 5.
@@ 2 The general model for axion double level crossings
where $\theta_{\mathcal{N}}$ and $\Theta_{i}$ are the $Z_{\mathcal{N}}$ axion and ALP angles, respectively, $f_{a_{\mathcal{N}}}$ and $f_{A_{i}}$ are the axion decay constants, $m_{a_{\mathcal{N}}}$ and $m_{A_{i}}$ are the axion masses, $n_{ij}$ are the domain wall numbers, and $\delta_{i}$ are the constant phases. The precise formula for the $Z_{\mathcal{N}}$ axion mass is detailed in appendix A, whereas the mass of ALPs, under the assumption of the simplest single-field scenario, remains constant across all instances. The constant phases $\delta_{i}$ are deemed to be zero, which is equivalent to seeking a separate resolution to the strong CP problem.
@@ 4 Redefining the light and heavy axion scenarios
Finally, we also show in figure 3 the first level crossing temperatures in both the light and heavy axion scenarios. The solid lines represent the temperatures $T_{\times_{i}}$ as functions of the value of $\mathcal{N}$}. The gray dashed lines represent the extreme scenarios where the values of $f_{A_{i}}$ and $f_{a_{\mathcal{N}}}$ differ greatly, $\rm i.e.$, $f_{A_{i}}\ll f_{a_{\mathcal{N}}}$ and $f_{A_{i}}\gg f_{a_{\mathcal{N}}}$, respectively. We find that, in the light axion scenario, variations in $T_{\times_{i}}$ exhibit sensitivity to large values of $\mathcal{N}$}, whereas, in contrast, in the heavy axion scenario, they are sensitive to small values of $\mathcal{N}$}. Notice that here we have selected some typical values of $f_{A_{i}}$ for illustration, particularly in the heavy axion scenario, where we have chosen three sets of $f_{A_{i}}$ such that they can be greater than, equal to, and less than $f_{a_{\mathcal{N}}}$, respectively.

### v2

@@
Hai-Jun Lia,b,c, Wei Chaod,e, Huai-Ke Guoc,f, and Yu-Feng Zhoub,c,f,g
aDepartment of Physics, Yunnan University, Kunming 650091, China
bInstitute of Theoretical Physics, Chinese Academy of Sciences, Beijing 100190, China
cInternational Centre for Theoretical Physics Asia-Pacific, Beijing 100190, China
dCenter for Advanced Quantum Studies, School of Physics and Astronomy, Beijing Normal University, Beijing 100875, China
eKey Laboratory of Multi-scale Spin Physics, Ministry of Education, Beijing Normal University, Beijing 100875, China
fSchool of Physical Sciences, University of Chinese Academy of Sciences, Beijing 100049, China
gSchool of Fundamental Physics and Mathematical Sciences, Hangzhou Institute for Advanced Study, UCAS, Hangzhou 310024, China
@@
In this work, we present the refinement of axion double level crossings within the context of multi-axion mass mixing, specifically focusing on cases where the number of axions exceeds two. Our investigation reveals that double level crossings are a common phenomenon in the mass mixing of the $Z_{\mathcal{N}}$ axion and axion-like particles. Physically, these double level crossings involve a first level crossing at high temperatures followed by a second level crossing induced by the $Z_{\mathcal{N}}$ axion mass transition at $T_{\rm QCD}$. We introduce the general model for double level crossings, along with several toy examples, and redefine the light and heavy axion scenarios. In the light axion scenario, double level crossings can occur multiple times in the large $\mathcal{N}$} limit. However, excessively large values of $\mathcal{N}$} may also prevent the occurrence of double level crossings. Conversely, in the heavy axion scenario, excessively small $\mathcal{N}$} may similarly prevent their occurrence. Our findings also have some intriguing implications for axion cosmology.
@@ 1 Introduction
Within such a theoretical framework of the axiverse, it is natural to consider the cosmological evolution of multiple axions, a subject that has garnered considerable attention in recent times Hill:1988bu ; Daido:2015cba ; Cyncynates:2023esj ; Li:2023uvt ; Li:2025cep . When considering non-zero mass mixing between axions, a significant evolution known as level crossing can occur during the QCD phase transition. Typically, level crossing precedes the critical temperature of the QCD phase transition, denoted as $T_{\rm QCD}$. However, as emphasized in ref. Li:2023uvt , this rule is not absolute; in scenarios featuring a $Z_{\mathcal{N}}$ axion, [^1] In the $Z_{\mathcal{N}}$ axion scenario Hook:2018jle ; DiLuzio:2021pxd , $\mathcal{N}$} mirror and degenerate worlds, nonlinearly realized through the axion field under a $Z_{\mathcal{N}}$ symmetry, can potentially coexist in Nature. Notably, one of these worlds corresponds to the SM world. To address the strong CP problem in this context, it is imperative that $\mathcal{N}$} be an odd integer satisfying ${\mathcal{N}}\geqslant 3$. The $Z_{\mathcal{N}}$ axion can also serve as a candidate for DM DiLuzio:2021gos . a second level crossing might occur precisely at $T_{\rm QCD}$, a phenomenon termed double level crossings (see also ref. Li:2024kdy for another illustration). Axion level crossing can induce the adiabatic or non-adiabatic transitions in axion energy density, similar to the MSW effect Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj in neutrino oscillations, and has some profound cosmological implications Kitajima:2014xla ; Daido:2015bva ; Ho:2018qur ; Cyncynates:2021xzw ; Cyncynates:2022wlq ; Li:2023xkn ; Li:2024psa ; Murai:2024nsp ; Li:2025uwq .
@@
The paper is structured as follows. In section 2, we present the general model for axion double level crossings. In section 3, we discuss several toy examples in the light axion scenario. Section 4 redefines the light and heavy axion scenarios. Section 5 briefly discusses the cosmological implications. Finally, section 6 concludes the paper.
@@ 2 The general model for axion double level crossings
where $\theta_{\mathcal{N}}$ and $\Theta_{i}$ are the $Z_{\mathcal{N}}$ axion and ALP angles, respectively, $f_{a_{\mathcal{N}}}$ and $f_{A_{i}}$ the axion decay constants, $m_{a_{\mathcal{N}}}$ and $m_{A_{i}}$ the axion masses, $n_{ij}$ the domain wall numbers, and $\delta_{i}$ the constant phases. The $Z_{\mathcal{N}}$ axion mass is detailed in appendix A, while the ALP mass remains constant in the simplest single-field scenario. The constant phases $\delta_{i}$ are deemed to be zero, which is equivalent to seeking a separate resolution to the strong CP problem. Since this assumption does not alter the evolution of the axion fields during mixing, it has no impact on the final axion production in our scenario Li:2023xkn .
@@ 4 Redefining the light and heavy axion scenarios
Finally, we also show in figure 3 the first level crossing temperatures in both the light and heavy axion scenarios. The solid lines represent the temperatures $T_{\times_{i}}$ as functions of the value of $\mathcal{N}$}. The gray dashed lines represent the extreme scenarios where the values of $f_{A_{i}}$ and $f_{a_{\mathcal{N}}}$ differ greatly, $\rm i.e.$, $f_{A_{i}}\ll f_{a_{\mathcal{N}}}$ and $f_{A_{i}}\gg f_{a_{\mathcal{N}}}$, respectively. We find that, in the light axion scenario, variations in $T_{\times_{i}}$ exhibit sensitivity to large values of $\mathcal{N}$}, whereas, in contrast, in the heavy axion scenario, they are sensitive to small values of $\mathcal{N}$}. Notice that here we have selected some typical values of $f_{A_{i}}$ for illustration, particularly in the heavy axion scenario, where we have chosen three sets of $f_{A_{i}}$ such that they can be greater than, equal to, and less than $f_{a_{\mathcal{N}}}$, respectively. Therefore, in the light axion scenario, a larger value $\mathcal{N}$} corresponds to a higher first level crossing temperature. When this temperature deviates significantly from that indicated by the gray dashed line, we consider the first level crossing to be compromised, thereby affecting the realization of double level crossings. Conversely, in the heavy axion scenario, the first level crossing can be compromised for smaller values of $\mathcal{N}$}, which similarly hinders the realization of double level crossings.
5 A brief discussion on cosmological implications
In this section, we briefly discuss the cosmological implications of axion double level crossings presented in this work.
Axion level crossing carries profound cosmological implications, encompassing axion relic density, isocurvature fluctuations, dark energy, domain walls, gravitational waves, and primordial black holes, as detailed in ref. Li:2025cep . Similarly, the axion double level crossings investigated here entail analogous cosmological consequences. While a comprehensive exploration of these effects is beyond the scope of this work and remains a promising avenue for future research, we briefly highlight below the novel insights our study offers regarding these phenomena.
Of these cosmological consequences, the modification of the axion relic density stands out as the most direct and crucial, serving as the foundation upon which the other effects depend. In this work, the parameter $\mathcal{N}$} plays a pivotal role. On the one hand, as previously discussed, the light and heavy axion scenarios impose upper and lower bounds on $\mathcal{N}$}, respectively. On the other hand, the $Z_{\mathcal{N}}$ axion model and its corresponding DM solution inherently constrain $\mathcal{N}$}, a restriction manifested in the $\{m_{a_{\mathcal{N}}},1/f_{a_{\mathcal{N}}}\}$ parameter plane. Consequently, we can constrain $\mathcal{N}$} by evaluating the degree of deviation from the ideal first level crossing, thereby narrowing the viable parameter space for the axion mass $m_{a_{\mathcal{N}}}$ and decay constant $f_{a_{\mathcal{N}}}$. Furthermore, here we consider the mixing of multiple ALPs with the $Z_{\mathcal{N}}$ axion. Given that ALP masses must be non-degenerate and exhibit a hierarchical structure, the number of ALPs participating in the double level crossings provides an additional constraint on $\mathcal{N}$}. Conversely, this limits the ALP parameter space, particularly the decay constants $f_{A_{i}}$ of ultra-light ALPs in the heavy axion scenario, where $f_{A_{i}}>f_{a_{\mathcal{N}}}/{\mathcal{N}}$. Notably, these ALPs fall within the sensitivity reach of future axion detection experiments, such as CASPEr-electric Budker:2013hfa and proposals based on the piezoaxionic effect Arvanitaki:2021wjk . Numerous additional searches for ALP couplings — specifically to photons, electrons, and nucleons — are also ongoing, though they are not discussed further here. Finally, while it is possible to directly calculate the relic density of the mixed $Z_{\mathcal{N}}$ axion or ALPs via the misalignment mechanism for a given model, such calculations are of secondary importance in this context due to the large number of model parameters, many of which remain undetermined.